OpenStudy (anonymous):

Use a 1-variable equation with rational coefficients to represent a situation involving two operations

OpenStudy (anonymous):

-3/4b \[\times\] 5 \[-\] 10 = 44

OpenStudy (anonymous):

A company needs to make 31.4 km of cable for a customer. The machine making the cable can create 5.4 km of cable per hour. While the company is working on the cable, the customer adds 5 km to the order. Which equation could be used to find how much time it will take to make all the cable? [A] 31.4 + 5 = 5.4x [B] 31.4x + 5 = 5.4 [C] 31.4 - 5.4x = 5 [D] 31.4x -5 = 5.4

OpenStudy (anonymous):

A because the original order was 31.4 km of cable, then the customer added 5 more km of cable. But the machine can still only make 5.4 km of cable and hour and so you add the x to the 5.4 to find out how many hours it will take

OpenStudy (anonymous):

Okay, makes sense! Thanks!

OpenStudy (anonymous):

Np. :)